# Comprehensive Review of Deep Unfolding Techniques for Next-Generation Wireless Communication Systems

### v2

@@ IV-A Massive MIMO
MIMO transmission uses multiple antennas at both the transmitter and receiver. It improves spectral and energy efficiency by utilizing a larger spatial domain. Adding more antennas (tens or hundreds) introduces the concept of massive MIMO [3]. However, detecting multiple symbols becomes more complex due to large-scale inter-symbol interference. Designing a low-complexity, high-performance detector for massive MIMO remains a challenge. For a MIMO communication system with $n_{\text{t}}$ transmit and $n_{\text{r}}$ receive antennas, the received signal can be given as:
$\mathbf{y}=\mathbf{Hx}+\mathbf{n}$ (1)
where $\mathbf{x}\in\mathbb{C}^{n_{\text{t}}}$ is the transmitted baseband signal vector, $\mathbf{H}\in\mathbb{C}^{n_{\text{r}}\times n_{\text{t}}}$ is the channel matrix and $\mathbf{y}\in\mathbb{C}^{n_{\text{r}}}$ is the received baseband signal vector. The detector extracts the desired transmitted information $\mathbf{x}$ from the received signal vector $\mathbf{y}$ with given $\mathbf{H}$ and the Gaussian noise vector $\mathbf{n}$.
The maximum a posteriori (MAP) and maximum likelihood (ML) estimations are Bayesian-based detection methods. These are the most commonly used classical detection techniques. The MAP rule applies when the transmitted signals have unequal a priori probabilities. The ML rule is used when these probabilities are equal [124]. ML minimizes the average error probability, providing an optimal solution. The ML detection rule is given as:
$\mathbf{\hat{x}_{\text{ML}}}=\text{arg}\min_{\mathbf{x}\in\mathcal{A}^{n_{t}}}||\mathbf{y}-\mathbf{Hx}||^{2}.$ (2)
This rule minimizes the Euclidean distance between $\bf{y}$ and $\bf{Hx}$. The minimization is performed over all possible combinations of the transmitted symbol vector $\bf{x}$ from the modulation alphabet $\mathcal{A}$. Since ML performs an exhaustive search and examines all possible signals, the computational complexity increases exponentially as $n_{\text{t}}$ increases.
MIMO detection can be linear or non-linear. Common linear detectors include matched filter (MF), zero-forcing (ZF), and minimum mean squared error (MMSE) [125]. These detectors require matrix inversion. Lattice reduction techniques can improve the performance of ZF and MMSE. Non-linear detectors, such as those using successive interference cancellation (SIC), perform interference estimation and removal in multiple stages [3], [126]. SIC and parallel interference cancellation are well-known techniques for this. In SIC, a detector like MF, ZF, or MMSE is used to detect the transmitted symbol in a data stream, leading to MF-SIC, ZF-SIC, or MMSE-SIC schemes. The data streams are arranged based on the received SNR. After detecting the symbol and estimating the channel coefficients, the interference caused by the detected symbol is estimated. This interference is then removed from the received signal, which is used to detect the next data stream in terms of received SNR. This process repeats until all symbols are detected. Sphere decoding is another non-linear detector [127]. It is an approximate ML detection method, where detection happens in a subspace of the modulation alphabet. The method searches within a bounded hypersphere around the received lattice point, using a radius for iterative symbol searching. The logarithmic likelihood ratio (LLR) is computed based on the symbol candidate found. This search region is smaller than the exhaustive ML search [128, 129].
The channel matrix for a MIMO fading channel can be modeled as a graph. The MIMO detection problem is framed as an inference task to calculate the posterior probability of system variables. Low-complexity algorithms are designed to compute these probabilities efficiently. The graph model combines probability and graph theory, showing the relationships between different nodes. Common graphical models include Bayesian belief networks, Markov random fields, and factor graphs [3]. Message-passing algorithms, like belief propagation (BP), are crucial for representing the MIMO system’s graph. Probability messages are exchanged between nodes representing transmitted symbols and received signals along the graph edges. The system is often modeled as a factor graph, where performance depends on the graph structure and algorithm convergence. Iterative algorithms, like message-passing detectors (MPD) with BP, approximate message passing (AMP), and channel hardening-exploiting message passing, provide solutions with lower complexity. Expectation propagation (EP) extends BP to handle larger constellations and higher-dimensional MIMO systems. EP approximates the posterior distribution of transmitted symbols iteratively, enabling efficient detection in polynomial time [130, 131].
The conventional detectors (linear and non-linear) involve computationally intensive matrix inversions or factorizations (like singular value decomposition or QR decomposition). Although these detectors yield nearly optimal solutions, they come at the cost of high complexity and high latency. Hence, with the aid of DL techniques, a deep unfolding approach has been adopted in literature to achieve low complexity and low latency.
PGD is an iterative optimization algorithm. Each step includes a gradient descent update to minimize a cost function, followed by a projection step to enforce constraints like sparsity or feasibility. In MIMO detection, PGD minimizes the error between the predicted and actual received signals while projecting the solution onto a feasible set, such as valid constellation points [132]. To achieve ML optimization, the PGD algorithm is unfolded into the DetNet architecture for MIMO detection [49, 133]. The gradient step size is learned across layers. The authors claimed that this approach reduces computational complexity compared to the sphere decoder and is robust to ill-conditioned channels. By leveraging the benefits of graphical models, He et al. proposed model-driven DL networks, OAMP-Net [10] and OAMP-Net2 [11], for MIMO detection with imperfect channel state information (CSI). OAMP-Net2 outperforms OAMP-Net by adapting to varying channel conditions. The orthogonal AMP (OAMP) algorithm is unfolded, and the LMMSE channel estimator is used to design the receiver. Learnable parameters, such as the step size for updating the prior mean and variance in the MMSE estimator, are trained using SGD. These parameters in OAMP-Net2 help compensate for channel estimation errors, improving performance compared to DetNet [49]. In [134], the authors introduced MMNet, a deep unfolded detector designed for a wide range of realistic channel matrices, including ill-conditioned and non-i.i.d. scenarios. MMNet is based on an iterative soft-thresholding algorithm, alternating between a linear detector and a non-linear denoising step. It balances architectural flexibility and computational complexity, outperforming DetNet [49] and OAMP-Net2 [11]. Overloaded massive MIMO systems, where the number of transmit antennas exceeds the number of receive antennas, require detectors with low complexity and good performance. In [61], the trainable projected gradient (TPG) algorithm was introduced for these systems. TPG uses a DL-aided iterative PGD method with trainable parameters, such as step size, optimized via backpropagation and SGD. TPG achieves lower computational cost and fewer trainable parameters compared to DetNet [49], OAMP-Net2 [11], and others. Additionally, a trainable detector based on the Hubbard–Stratonovich transformation is discussed in [135]. Takabe et al. avoided matrix inversion in this approach. Compared to the TPG detector [61] and OAMP-Net2 [10], the proposed network achieves faster training and lower execution cost.
In [136], the concrete MAP method is introduced, based on a continuous relaxation of the probability mass function for discrete random variables, as derived in [137]. This approach solves the MAP detection problem using continuous optimization instead of exhaustive search, approximating the MAP solution for any probabilistic nonlinear model. By unfolding the gradient descent algorithm, the detector provides soft outputs (detection probabilities), enabling better decoding compared to DetNet [49], OAMP-Net2 [10], and MMNet [134]. Yun et al. in [138] enhanced the unfolded PGD design by adding momentum components to the iterative hidden layers. Momentum helps guide parameter updates in the right direction and reduces oscillations during PGD optimization [139]. This modification improves convergence speed and overall performance, outperforming DetNet [49] and OAMP-Net2 [10, 11], while maintaining generality.
In wireless communication, physical layer signal processing involves continuous-time analog waveforms, which must be quantized for hardware implementation. Low-resolution one-bit analog-to-digital converters are commonly used in massive MIMO systems due to their high sampling rates, low cost, and low energy consumption [140]. To recover high-dimensional signals from noisy, low-resolution measurements, Khobahi et al. [141] developed a low-resolution detection network. This network extracts signals from noisy one-bit measurements without requiring prior knowledge of the channel matrix. It is based on unfolding the PGD method for symbol detection using the ML estimator, eliminating the need for CSI. The network outperforms traditional ML estimators [142] and DL networks like DeepSIC [143], even with small datasets of about 500 samples. In [59], a generalized PGD algorithm was proposed, eliminating the need for separate loss functions for projection and gradient steps. Instead of alternating after each step, projection is applied after several gradient steps, achieving a better balance. By unfolding this approach, the authors created a self-correcting auto detector that enhances performance [59].
Homotopy optimization transforms the original problem into a different one using a homotopy map. This method solves a series of simpler problems step by step, gradually progressing to the harder ones [144]. In MIMO detection, non-convex algorithms often struggle with the ML detection problem due to bad local minima [50]. By using a homotopy map to transition from easier to harder problems, detection performance is significantly improved. The study in [50] focuses on ML MIMO detection with one-bit quantized observations and binary symbol constellations. Using homotopy optimization, the authors apply a non-convex continuous reformulation of binary optimization [145] to address the detection problem. Their algorithm is designed to run efficiently and adapt to different channels. Building on this work, the authors introduced a proximal gradient-based homotopy method in [53]. This approach connects the structure of DetNet [49] with homotopy optimization for improved MIMO detection.
GNNs, as structured DNNs, excel in adaptability and flexibility. In MIMO detection, the graph structure of wireless networks makes GNNs particularly effective. Tan et al. [55] unfold MPD iterations to update messages at symbol nodes using prior and posterior LLR. Learnable parameters, such as prior probabilities and correction factors (damping, rescaling, offset), are optimized via mini-batch SGD to minimize the loss function. A greedy search determines the number of layers to balance performance and complexity. In [60], a low-complexity AMP-GNN model is proposed for massive MIMO detection. It outperforms OAMP-Net2 [10] by avoiding matrix inversion. For single-carrier FDMA systems, common in 5G to reduce peak-to-average power ratios, message-passing algorithms are computationally demanding. To address this, the unfolded network in [146] simplifies generalized AMP iterations [147], maintaining high performance even with channel mismatches. To enhance EP-based detection, [51] optimizes parameter selection using a modified moment-matching scheme, showing superior performance in high modulation cases compared to OAMP-Net2. In [52], approximate EP from [148] is unfolded, achieving better detection performance and robust convergence.
Liao et al. [44] incorporated interference cancellation into DNN layers using iterative embedding. Inspired by [149], their sequential detector trains residual error coefficients with MSE loss, reducing complexity by avoiding matrix inversion and relying on matrix addition and multiplication. Compared to DetNet [49, 133], which requires 90 layers for convergence, their model achieves convergence with just 8 layers, offering significant computational simplicity. Massive MIMO systems improve spectral efficiency (SE), coverage, and range but demand higher computational power. LcgNet, a model-driven DL architecture for massive MIMO detection, unfolds conjugate gradient descent iterations and learns step sizes through offline training [54]. Its quantized version, QLcgNet, minimizes memory usage with negligible performance degradation. LcgNet outperforms DetNet in terms of symbol error rate, computational complexity, and memory cost, making it a robust solution for massive MIMO systems.
The quantized version, QLcgNetV, reduces memory requirements while maintaining high detection performance. To enhance LcgNet, Olutayo et al. [150] incorporated a preconditioner during training, designing a preconditioned LcgNet. Preconditioning improves the condition number of a matrix [151], optimizing the receiver filter’s spectrum and minimizing eigenvalue spread. This enhancement significantly boosts the convergence speed compared to the original LcgNet [54]. Additionally, the dynamic conjugate gradient network utilizes self-supervised learning with forward error correction codes, adapting effectively to non-stationary, time-varying channels.
Sphere decoding achieves ML optimality but suffers from high computational complexity in massive MIMO systems. To address this, Nguyen et al. [123] proposed a fast DL-aided sphere decoding algorithm inspired by [152]. The DNN generates reliable initial candidates for sphere decoding, prioritizing candidates closer to the DNN’s output for testing. This approach accelerates sphere shrinkage, reducing computational cost. The sparse network architecture eliminates matrix inversion, relying only on element-wise matrix multiplications. The proposed method performs better in highly correlated channels than DetNet [49, 133], OAMP-Net2 [10, 11], MMNet [134], and LcgNet [54]. A MIMO detector based on annealed Langevin dynamics is introduced in [57]. It replaces the traditional two-step iterative framework (continuous-space optimization followed by projection) with a single-step approximation of the MAP estimator by sampling from Langevin dynamics. This approach shows superior performance in correlated channels compared to DetNet [49], OAMP-Net2 [11], and MMNet [134]. To mitigate the computational costs of matrix inversion in classical detectors like ZF and MMSE, Berra et al. [58] proposed unfolding accelerated Chebyshev successive over-relaxation and over-relaxation algorithms. This network requires fewer trainable parameters and features a fast, stable training scheme. The unfolded detector achieves better performance compared to DetNet [133], OAMP-Net2 [10], TPG-detector [61], MEPD [51], EP-Net [46], and other methods like [53, 44].
To avoid the computational burden incurred due to matrix inversion, the authors in [56] have proposed a sparse refinement architecture for symbol detection in uplink massive MIMO. The authors have used deep unfolding for various crucial components in matrix inversion, significantly reducing the detection complexity. For detection, the alternating direction method of multipliers (ADMM) algorithm is utilized in [153]. ADMM decomposes a complex optimization problem into simpler subproblems, which are solved iteratively. The algorithm combines the principles of dual ascent and the method of multipliers. Each iteration involves solving the subproblems for the primal variables in an alternating manner, followed by updating the dual variables (Lagrange multipliers) to enforce consistency with the original problem’s constraints [154]. This iterative process ensures that the subproblem solutions converge to the global solution of the main optimization problem. The authors in [155] developed an unfolded ADMM-based network for MIMO detection. They employed an inexact ADMM update [154, 156] to reduce complexity while achieving near-optimal performance. Compared to DetNet [49, 133], the proposed network requires fewer trainable parameters, resulting in improved performance in terms of symbol error.
IV-B Modern Multicarrier Modulation
An OFDM-based MIMO detector faces many challenges. The increased number of antennas at the transceiver exhibits enhanced spatial multiplexing but at the cost of increased inter-symbol interference (ISI). In standard OFDM multicarrier modulation, the cyclic prefix is inserted between OFDM blocks to mitigate ISI. Zhou et al. in [90], have unfolded a conjugate gradient method based OAMP detector for the OFDM-MIMO system. To address the issues faced in detecting coarsely quantized or low-resolution signals in OFDM-based MIMO systems, Shao et al. [91] have unfolded the expectation maximization (EM) algorithm. By unfolding the likelihood ascent search algorithm, Ullah et al. [92] are able to jointly detect and estimate the soft output in MIMO OFDM systems.
Li et al. [46] proposed a deep unfolding network for OTFS detection by converting the iterative EP algorithm into structured network layers. The core idea is to approximate the conditional probability distributions $p(\bf{y}|\bf{x},\bf{H})$ and $p(\bf{x})$ where $p(\cdot)$ represents the probability, and $\bf{y},\bf{x},\bf{H}$ denote the received signal vector, transmitted signal vector, and channel matrix, respectively. The trainable parameters in this network include the step size, damping factor, and offset values for the mean ($m$) and variance ($v$) of the probability distribution. During offline training, the number of layers in the unfolding network is selected to balance performance and computational complexity. To enhance convergence and prevent the network from being trapped in local optima, the Adam optimizer with warm restarts is employed. This technique adaptively adjusts the step sizes, improving both training stability and performance. In [101], an unfolded OAMP-based detector for OTFS is introduced to handle channel estimation errors. This detector is trained using imperfect CSI and optimizes three trainable parameters: step sizes for updating the linear estimate, nonlinear estimate, and noise variance.
IV-C Next Generation Multiple Access
A cellular network must simultaneously support a growing number of active subscribers within a limited time-frequency spectrum [157]. The current trend in wireless communication is to increase the signal bandwidth to achieve higher transmission rates. However, increasing bandwidth results in shorter symbol periods due to the inverse relationship between bandwidth and symbol period. Consequently, in a multi-path fading channel, the reduced symbol period exacerbates ISI. Multi-carrier modulation techniques like OFDM provide an effective solution to mitigate ISI by dividing a wideband signal into a set of orthogonal narrowband subcarriers [158]. This approach not only minimizes interference but also enhances robustness in multipath environments.
Multiple access techniques allow multiple users to share the same communication channel and resources, such as time, frequency, or code, to transmit data efficiently. Traditional multiple access schemes achieve this by dividing signaling dimensions orthogonally across time, frequency, or code [159]. This orthogonality enables low-complexity multi-user detection at the receiver. However, orthogonal multiple access cannot achieve the desired sum-rate capacity for a multiuser wireless system. To address this, NOMA has been introduced, allowing more users to be served than the number of orthogonal resource units by leveraging non-orthogonal resource sharing [157]. NOMA is categorized into power-domain NOMA and code-domain NOMA [160]. In power-domain NOMA, the transmitter superimposes information symbols into a single waveform using superposition coding. At the receiver, SIC is employed to iteratively decode signals until the desired signal is recovered.
A deep unfolded detector, named the learned preconditioned conjugate gradient descent network with SIC, is proposed for downlink MIMO-NOMA systems in [105]. To achieve a robust model with low dataset requirements, the algorithm incorporates a data transfer learning strategy. This approach adapts the model efficiently to variations in environmental factors, such as channel models, modulation schemes, or power allocation, ensuring consistent performance across diverse scenarios.
In sparse spread CDMA, each user is assigned spreading codes with only a few non-zero elements. This sparsity minimizes overlap and enables multiple users to transmit simultaneously on the same resources with minimal interference. In sparse CDMA, the spreading matrix contains sparse non-zero elements, as initially studied in [161]. Each user’s transmitted symbols are encoded using unique sparse signature sequences, which facilitate decoding at the BS [162]. Takabe et al. proposed a complex sparse trainable projected gradient detector (C-STPG) for SCDMA networks in [106]. The detector builds upon the PGD framework, delivering significantly improved detection performance compared to conventional LMMSE detectors. Simulation results demonstrate reduced computational complexity during execution and minimal training costs due to the limited number of trainable parameters. The authors also introduced gradual sparsification—a joint learning technique that designs sparse signature sequences using deep unfolding. This technique optimizes the weights of the signature matrix while reducing error during training through MSE loss. The trainable parameters include the step size for gradient descent ($\gamma$) and a softness parameter ($\alpha$) to adjust the strength of the soft projection.
IV-D Orbital Angular Momentum
[…]

### v3

@@ IV-A Massive MIMO
MMSE-based channel estimation is commonly used in MIMO systems to estimate the channel response in noisy environments. In massive MIMO systems with frequency division duplexing (FDD), reducing CSI feedback overhead is challenging. A method in [99] combines DL and superposition coding to address this. Downlink CSI is encoded, superimposed on uplink user data, and sent to the BS. A multi-task NN unfolds the MMSE process to reduce interference and reconstruct the CSI efficiently.
ISTA is commonly used for sparse channel estimation, enabling efficient CSI feedback. Each ISTA iteration includes a linear operation followed by a nonlinear soft-thresholding step, similar to the rectified linear unit (ReLU) activation function [20]. In [100], Chen et al. proposed a trainable DL architecture for sparse recovery problems using an adaptive depth approach. Instead of a fixed number of layers, the method dynamically adjusts the depth based on the problem’s complexity. Highly sparse channels use fewer layers, while more complex channels require additional layers. This adaptive approach improves accuracy and lowers computational costs compared to classical methods and fixed-depth networks. Results in [100] demonstrate that the method reduces reconstruction error and speeds up channel estimation, making it more efficient. A scenario-adaptive CSI compression framework is proposed in [101], requiring minimal channel measurements for training. The authors introduced a CSI feedback architecture designed for efficient encoder updates in dynamic wireless environments, which is essential for downlink precoding in massive MIMO systems. The encoding network focuses on improving the robustness and efficiency of CSI compression. It employs a deep unfolding-based feedback network to iteratively refine the feedback process and enhance feature extraction. Similarly, the decoding network uses a deep unfolding structure inspired by ISTA. This iterative approach improves CSI recovery accuracy, making the process more efficient and adaptable to changing conditions.
In massive MIMO systems, where the channel matrix is often sparse, the majorization-minimization (MM) approach is effective for addressing sparse channel estimation [102]. In [103], the authors address the challenge of downlink channel estimation in massive MIMO. A two-timescale joint uplink/downlink dictionary learning and channel estimation algorithm is proposed. The short-term subproblems focuses on estimating sparse channel vectors for each timeslot. The MM algorithm is unfolded, and the sparse channel vectors are treated as trainable parameters. In contrast, the long-term subproblem involves updating the dictionary to adapt to dynamic environments for which a constrained stochastic successive convex approximation (SSCA) framework is employed without deep unfolding. Simulations results in [103] demonstrate that the proposed approach reduces pilot overhead and outperforms existing methods in both static and dynamic environments, particularly in environments with mobile users.
The matching pursuit algorithm is a greedy iterative method often applied in sparse channel scenarios like massive MIMO. It identifies channel components (basis functions) by matching them to the residual, which is the difference between the observed signal and the signal generated by the current channel estimate. Yassine et al. [104] introduced a deep unfolded algorithm, called mpNet, based on the classical matching pursuit algorithm. This network estimates the multiuser massive MIMO channel in an unsupervised manner, initializing with a dictionary of imperfect steering vectors. The system operates in time division duplex mode to leverage channel reciprocity. The network dynamically adjusts the number of layers based on incoming data, ensuring efficient performance in realistic environments.
The basis pursuit algorithm is a widely used method for sparse signal recovery, particularly in compressed sensing. It identifies the sparsest solution to an underdetermined system by minimizing the $\ell_{1}$-norm of the signal. This approach is crucial in scenarios where data is acquired under constraints, and the signal is assumed to be sparse in certain domains. In [105], the authors observed that existing sparse channel estimation schemes for downlink CSI often deliver suboptimal reconstruction due to the use of random measurement matrices. To address this, they proposed deep unfolding-based autoencoders, that optimize the measurement matrix by unfolding the basis pursuit algorithm. This customized design enhances reconstruction performance in sparse channel estimation tasks.
Sparse Bayesian learning is another effective approach for channel recovery [106]. It estimates the non-zero elements of the channel matrix while utilizing prior knowledge of sparsity. Hu et al. [107] introduced a deep unfolding framework called deep deterministic policy gradient (DDPG) for MU-MIMO systems. DDPG, a variant of deep Q-networks, is designed to handle problems in continuous action spaces [108]. Unlike conventional deep unfolding NNs with fixed depth, the authors leveraged DDPG to dynamically adjust the network depth based on the complexity of the problem. This adaptive design enhances the efficiency and performance of sparse channel recovery in varying scenarios. The channel estimation problem is formulated using an off-grid sparse Bayesian learning model. The sparse Bayesian learning algorithm is unfolded with trainable parameters grid points and off-grid variables. Simulation results show that the proposed method significantly outperforms fixed-depth models, achieving a 30% reduction in the number of layers while improving normalized MSE (NMSE) performance.
IV-B mmWave Massive MIMO
mmWave massive MIMO operates in the high-frequency range of 30 GHz to 300 GHz, offering abundant spectral resources to alleviate the bandwidth crunch at sub-6 GHz frequencies [82]. However, mmWave signals suffer from severe propagation and penetration losses. To compensate for these losses, large-scale antenna arrays with tens to hundreds of elements are required to achieve high power gains. The short wavelength (1–10 mm) of mmWave signals allows compact integration of these arrays into small transceivers. These features introduce unique challenges in designing physical-layer transmission algorithms [109]. The BS, equipped with numerous antennas, must acquire downlink CSI. However, the requirement of large amount of feedback creates significant overhead, posing a challenge in channel estimation in mmWave massive MIMO systems.
Compressive sensing techniques are widely used to reduce feedback overhead by exploiting channel sparsity [110]. The limited scattering properties of mmWave propagation and the power-focusing capability of lens antenna arrays transform the spatial domain channel into a sparse beamspace representation. Consequently, the channel estimation problem becomes a sparse signal recovery task, often addressed through compressed sensing methods. However, performing beamspace channel estimation in massive MIMO systems with limited RF chains remains a challenging task.
Compressive sensing aided deep unfolding is explored in [111], where the authors apply Onsager correction for AMP as explained in [112]. The architecture of the unfolded model, learned AMP (LAMP) [111] is similar to unfolded ISTA in [20]. A learned denoising-based AMP network (LDAMP) is proposed in [113] to learn beamspace channel estimates. The authors demonstrate that the iterative sparse signal recovery algorithm with denoising achieves robustness in channel estimation particularly in scenarios with large antenna arrays and limited RF chains. Yi et al. in [114] introduced an unfolded network called learned AMP with deep residual learning (LampResNet). In this model, the AMP algorithm is unfolded to estimate the channel, and deep residual learning is employed to mitigate channel noise. NMSE analysis shows that the proposed network in [114] outperforms LAMP [111] and achieves performance comparable to LDAMP [113] in terms of estimation accuracy. However, exploiting channel sparsity does not always yield optimal estimation performance. To address this, the authors of [115] investigated the sparsity of the mmWave channel in the beamspace. They formulate the estimation problem as a sparse signal recovery task and improve convergence speed by unfolding the iterative PGD algorithm. In this approach, the step-size of the descent step is treated as a trainable parameter. In [116], an iterative trimmed-ridge regression algorithm is unfolded to improve beamspace channel estimation. The proposed algorithm involves parallel computation to achieve negligible latency. These algorithms were applied to both narrowband and wideband massive MIMO systems, to ensure versatility in channel estimation.
IV-C Modern Multicarrier Modulation
Yiyun et al. [117] proposed a CSI recovery network for FDD downlink massive MIMO-OFDM systems. ISTA is unfolded and the network consists of compression and reconstruction modules for CSI compression and feedback. Building on ISTA, Hu et al. [118] introduced an unfolded algorithm called the learnable optimization and regularization algorithm (LORA) to enhance the precision of CSI feedback. This approach also unfolds ISTA, and the regularization term is learned through end-to-end training. The model captures CSI characteristics and adapts to different levels of channel estimation error. Furthermore, to improve CSI feedback accuracy and efficiency, Yangyang et al. [119] proposed an unfolded ISTA network. Their study showed that the real and imaginary components of the channel differential terms, i.e, the difference between CSIs at adjacent time slots, exhibit similar correlation patterns in both the angular and delay domains. This property enables a shared model to process both components, thereby reducing network complexity and enhancing the efficiency of the CSI encoder. Each of the networks proposed in [117, 118, 119] enhance CSI feedback performance for FDD massive MIMO-OFDM system. In [33], Guo et al. proposed a two-stage low-rank CSI feedback scheme for a low-rank OFDM mmWave channel. In massive MIMO systems operating in FDD mode, CSI is estimated by unfolding the fast ISTA. Fast ISTA is an improved version of the classical ISTA due to the introduction of momentum term that accelerates the convergence rate. The network, FISTA-Net, includes trainable step-size and thresholds to achieve accurate reconstruction performance. In [120], the authors explore quantized compressed CSI at the user equipment before feedback for massive MIMO-OFDM system. They propose a deep unfolding-based bit-level CSI feedback network, TiLISTA, to mitigate the effects of accumulated quantization errors. Building upon ISTA, they used tied learned ISTA as discussed in [121]. Since the trainable parameters are shared across all layers, the number of learnable parameters are reduced, offering a faster and more accurate solution to the signal reduction problem. FISTA-Net [33] focuses on faster convergence while TiLISTA [120] targets robustness to quantization errors.
In [122], an unfolded sparse Bayesian learning algorithm is derived to capture channel sparsity in a mmWave massive MIMO-OFDM system with hybrid beamforming (HBF). The proposed network achieved a lower NMSE of $0.033$ outperforming the classical sparse Bayesian learning algorithm which recorded an NMSE of $0.1203$. Jianqiao et al. [123] proposed an unfolded inverse-free variational Bayesian learning framework to address the high computational complexity of matrix inversion in mmWave MIMO-OFDM systems. The sparse channel recovery performance is enhanced by learning the a priori parameters of the sparsely represented dictionary for channel estimation. In [124], based on approximate Bayesian inference and an extension of expectation-consistent approximation, the authors proposed a generalized expectation consistent algorithm, an iterative message-passing approach for sparse channel estimation. In a lens-based beamspace massive MIMO-OFDM system, this approach compensates for large channel attenuation in the mmWave band. The authors developed a deep unfolded unsupervised learning network called the learned denoising-based generalized expectation consistent signal recovery network. The generalized expectation consistent signal recovery algorithm as discussed in [125] is unfolded. Both the networks in [123] and [124] provide accurate channel recovery estimates for mmWave channel with few RF chains. In [126], the authors leveraged angle-domain channel sparsity to reduce uplink pilot overhead during channel estimation for mmWave MIMO-OFDM systems. They propose an unfolded multiple-measurement-vectors-based AMP network. The phase shift network and channel estimator are jointly trained, thus integrating the channel sparsity for improved performance.
Gao et al. [127] proposed a network, ComNet, that unfolds an OFDM receiver into two subnetworks: channel estimation and signal detection. The LMMSE and ZF are unfolded for channel estimation and signal detection, respectively. The weights of the LMMSE and ZF matrices are trained layer-wise. The simulation results show that the BER performance is better compared to the classical LMMSE estimation. Furthermore, Wang et al. in [128] proposed an unfolded least squares (LS) estimation to enhance channel estimation for OTFS MIMO system under fractional Doppler condition. The network, CENet, is structured as parallel sub-networks, each responsible for estimating one row (delay tap) of the delay-time channel matrix. At BER= $3\times 10^{-3}$, CENet achieved $\sim 5$ dB SNR gain over LS estimation algorithm.
TABLE IV: Summary of typical deep unfolding models for channel estimation and their advantages and disadvantages.
Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
LDAMP [113] Denoising-based unfolded AMP model for sparse-beamspace channel estimation in mmWave massive MIMO. Outperforms classical denoising AMP method with small number of RF chains at the receiver. Advantages: Simple architecture and fast convergence (within $5$ layers), require a few RF chains. Disadvantages: Fixed parameters post-training restrict the model’s adaptability, performance may degrade in varying scenarios.
LampResNet [114] Combines unfolded AMP with deep residual learning to enhance robustness against noise. Achieves close performance to LDAMP [113] at high SNR regime ($25$ dB). Advantages: Resilient to noise, consistent NMSE improvement under SNR variations. Disadvantages: Rely on prior information of beamspace channel, not consistent with realistic channel models, and network may fail to converge.
FISTA-Net [33] Fast ISTA-based deep unfolding for low-rank CSI feedback in mmWave OFDM systems. Outperforms the classical FISTA with lower complexity. Advantages: Fast convergence ($20$ layers). Disadvantages: Requires retraining for major environment variations.
mpNet [104] Deep unfold model based on the matching pursuit method for multiuser massive MIMO systems. Outperforms classical LS estimator in terms of NMSE under varying SNR. Advantages: Avoids ground-truth channel data by using unsupervised online learning on noisy LS estimates. Disadvantages: Need prior knowledge of physical wave propagation model, as small uncertainty of system parameters lead to high estimation performance loss.
LORA [118] The deep unfolding model based on ISTA introduces a learnable regularization module to adaptively capture the CSI structure. Improved CSI reconstruction accuracy by $>2$ dB NMSE over TiLISTA [117]. Advantages: LORA exhibits strong denoising capability under channel estimation error. Disadvantages: The number of layers is critical: using fewer than four layers degrades reconstruction accuracy, while increasing the number beyond four leads to higher computational complexity with only marginal improvements in NMSE.
IV-D IRS
To mitigate the blockage effect and extend the connectivity range for mmWave communication, IRS is a key technology proposed to enhance both energy and SE. IRS consists of many reflecting elements that steer the propagation of electromagnetic waves by adjusting their reflection coefficients. However, to unlock the full potential of IRS, accurate channel estimation is required for both direct and cascaded links. This is a challenging task because IRS-based channel estimation involves estimating multiple channels simultaneously: the direct channels between the BS and each user, the channel between the BS and IRS elements, and the channel between IRS elements and users.
In [129], Jiguang et al. proposed a method for estimating the rank-deficient cascaded channel in an IRS-assisted single-input multiple-output system. They developed a deep unfolding network based on a gradient descent algorithm. The model structure is inspired by DetNet [49]. In massive MIMO, increasing the number of IRS elements and BS antennas enhances the system’s capacity. However, this leads to increased training overhead to obtain accurate channel estimates within the channel coherence time. The proposed network addresses this challenge by compensating for the computational complexity.
Awais et al. [130] developed two networks based on the MAP estimation criterion to address the challenges in IRS cascaded channel estimation. The first network is a deep denoising network designed for noise-dominated SNR regions, providing support at various noise levels. The second network is a deep unfolded half-quadratic splitting network that alternates between optimizing extrapolation and denoising for interference-dominated regions. As the SNR increases, the proposed network shows significant improvement in NMSE. Under severe channel variations, the network delivers a performance gain of 20% in the high SNR range.
Tsai et al. [131] proposed a two-stage learned AMP network with row compression to address the joint estimation problem of direct and cascaded channels in IRS-aided mmWave systems. By utilizing the deep unfolding technique, the network combines the advantages of compressive sensing and DL. This approach enables reliable channel estimation with reduced training, which is crucial for efficiently reconfiguring wireless environments under IRS reflection control.
IV-E IoT devices
Joint activity detection and channel estimation are critical in grant-free random access for massive machine type communication (mMTC), where numerous IoT devices access the network sporadically without scheduling. To enhance detection under imperfect CSI, Qiang et al. [31] proposed AMP-Net, an unfolded AMP based network. Each AMP iteration alternates between MMSE-based denoising and residual updates to minimize the MSE. The network uses four trainable scalar parameters to estimate activity probability, channel variance, regulate the denoiser, and control the state evolution. AMP-Net achieved accurate detection of the active device by extracting the CSI from the estimated device state matrix with a $4$ dB gain over classical AMP at SNR$=15$ dB. The authors in [132] proposed a grant-free access protocol for mmWave cell-free massive MIMO using a hybrid structure based on learned vector AMP, reducing reliance on prior channel knowledge and improving convergence over classical AMP. In [133], Ma et al. introduced a one-phase non-coherent scheme for mMTC. The data bits are embedded in pilot sequences to enable joint activity and data detection without channel estimation. They used unfolded AMP to leverage pilot correlation. In [134], an unfolded AMP with backpropagation was applied to an asynchronous random access scheme with data length diversity, addressing activity detection, channel estimation, and data recovery in mMTC.
[…]